OpenStudy (anonymous):

The line that passes through (-2, 4) and is parallel to x - 2y = 6

OpenStudy (anonymous):

parallel a = a'

OpenStudy (raden):

one of way : The line that passes through (p, q) and is parallel to ax + by = c is ax + by = ap + bq

OpenStudy (raden):

so, The line that passes through (-2, 4) and is parallel to x - 2y = 6 is x - 2y = -2 - 2(4) continue it ....